OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles?

Mehek (mehek14):

25/2=?

Mehek (mehek14):

first you have to find out how many miles is 1 inch

OpenStudy (igreen):

Set up a proportion: \(\dfrac{2}{25} = \dfrac{x}{60}\) Multiply 60 to both sides: \(x = \dfrac{60 \times 2}{25}\) Simplify

OpenStudy (anonymous):

4.8?

OpenStudy (igreen):

Yep!

OpenStudy (igreen):

That's our answer.

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

I need help in an other one

OpenStudy (anonymous):

What is the area, in square feet, of the shaded region?

OpenStudy (igreen):

One question per post, close this and open a new one.

OpenStudy (igreen):

Oh, and \(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):